if 0.650 solution of na2co3 will be use to precipitate ca+2 from solution, determin the amount of solution in grams needed to precipitate 5.00x102ml of 0.01120m ca+?(atomic mass:na=22.99 g/mol;c=12.01g/mol;0=16.00g/mol;ca=40.08g/mol;ans.91.3g

1
Expert's answer
2022-04-28T09:36:01-0400

"Ca^{2+}+ Na_2CO_3 \rightarrow CaCO_3 + 2Na^+"


At 5.00x102ml of 0.0112 m solution is "5x102x0.0112\/1000=0.005712 mole" of "Ca^{2+}"

From reaction is that "n(Na_2CO_3)=n(Ca^{2+})=0.005712"

"M(Na_2CO_3)=22.99x2+12.01++16x3=105.99 g\/mol"

So 0.005712 is "0.005712x105.99=0.6054 g" of "Na_2CO_3"

As solution concentration is 0.65, then it needs 0.6054/0.65=0.9314 g of solutuon.